if vector P + vector Q = vector R and if mode of vector P + mode of vector Q = mode of vector R then what is the angle between vector P and vector Q.
 

Grade:11

1 Answers

Arun
4 years ago
Let P=Q=R=AMagnitude of vectorR is|R|=(P^2+Q^2+2PQcosthita)where thita is angle between p and qA^2=A^2+A^2+2A^2costhitaCosthita=-1/2Thita=120*Hence angle between p and q is 120 degree